40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Sep 23 19:36:41 2009, rana, Update, PSL, RC temperature performance Untitled.pngUntitled.pnge.png
    Reply  Mon Nov 23 10:11:03 2009, steve, Update, PEM, long term temp fluctuation of the 40m lab 7ytemp.jpg
Message ID: 1995     Entry time: Wed Sep 23 19:36:41 2009     Reply to this: 2312
Author: rana 
Type: Update 
Category: PSL 
Subject: RC temperature performance 

This first plot shows the RC temperature channels' performance from 40 days ago, before we disabled the MINCO PID controller. Although RCTEMP is supposed to be the out of loop sensor, what we really care about is the cavity length and so I've plotted the SLOW. To get the SLOW on the same scale, I've multiplied the channel by 10 and then adjusted the offset to get it on the same scale.

 The second plot shows a period after that where there is no temperature control of the can at all. Same gain scaling has been applied to SLOW as above, so that instead of the usual 1 GHz/V this plot shows it in 0.1 GHz/V.

The third plot shows it after the new PID was setup.

Summary: Even though the PID loop has more gain, the true limit to the daily fluctuations in the cavity temperature and the laser frequency are due to the in-loop sensors measuring the wrong thing. i.e. the out-of-loop temperature is too different from the in-loop sensor. This can possibly be cured with better foam and better placement of the temperature sensors. Its possible that we're now just limited by the temperature gradients on the can.

Attachment 1: Untitled.png  131 kB  | Hide | Hide all
Untitled.png
Attachment 2: Untitled.png  15 kB  Uploaded Wed Sep 23 20:48:19 2009  | Hide | Hide all
Untitled.png
Attachment 3: e.png  17 kB  Uploaded Wed Sep 23 20:55:59 2009  | Hide | Hide all
e.png
ELOG V3.1.3-